# Training-Free Token Pruning via Zeroth-Order Gradient Estimation in Vision-Language Models

### v1

@@ 1 Introduction
Large Vision–Language Models (VLMs) (Bai et al., 2023; Liu et al., 2024b; Zhu et al., 2025) achieve strong multimodal understanding but at a substantial computational cost. A major contributor to this cost is the large number of visual tokens produced by modern vision backbones. For example, the vision encoder in LLaVA-1.5 (Liu et al., 2024a) generates up to 576 tokens for a single image, whereas the text side often contains only a few tokens, such as “Describe this image in a short sentence”. This imbalance leads to high inference latency and memory overhead (Chen et al., 2024a; Alvar et al., 2025). To address this issue, token pruning has emerged as a practical solution that selectively removes less informative visual tokens at inference time. Prior work has shown that aggressive token reduction can yield large speedups with only modest drops in accuracy (Zhang et al., 2025; Yang et al., 2025).
@@
Recent research on VLM token pruning has increasingly focused on training-free schemes. Unlike methods that require calibration data or fine-tuning (Hudson & Manning, 2019; Li et al., 2025; Lin et al., 2025), training-free approaches prune tokens directly at inference. These methods can be broadly categorized into two groups. (1) Attention-based approaches score tokens using attention magnitudes; however, attention often concentrates on background regions (Darcet et al., 2023) and tends to retain redundant tokens with overlapping content (Yang et al., 2025; Shang et al., 2024). For example, in an image of a laptop on a desk (Fig. 1), attention-based pruning may preserve many redundant tokens with overlapping content from the background, while ignoring tokens near the monitor that are critical for answering a question about objects on the monitor. (2) Diversity-based approaches (Alvar et al., 2025) select tokens by maximizing feature diversity, measuring pairwise distances between token embeddings. However, because these methods prioritize diversity uniformly across all tokens without explicitly considering task-relevant cues, they may discard tokens from visually salient regions. In the same laptop example, diversity-based pruning may select feature-diverse tokens, but not necessarily those near the monitor, which are crucial for reasoning about nearby objects.
Figure 1: Illustration of training-free VLM token pruning methods. (a) Attention-based methods select tokens using attention scores, but often retain redundant tokens. (b) Diversity-based methods select tokens with different features to maximize coverage but may lose tokens located in semantically relevant regions (e.g., around the monitor, highlighted in yellow). (c) Our method employs zeroth-order gradient estimation to quantify token sensitivity and integrates these scores into a diversity objective. (d) Accuracy comparison with LLaVA-NeXT-7B across 9 benchmarks, showing that ours outperforms both VisionZip (attention-based) and DivPrune (diversity-based) methods.
@@
Given these limitations, selecting a smaller subset of tokens while preserving visual information without degrading task performance remains a challenging problem. To tackle this, we investigate a novel metric for scoring each token, explicitly considering token sensitivity. Directly computing token sensitivities via gradients is costly, especially for VLM, motivating the use of zeroth-order gradient estimation (Nesterov & Spokoiny, 2017). This method quantifies how perturbations in input tokens influence model outputs using only forward passes, avoiding backpropagation. However, a naive application of zeroth-order estimation would require additional forward passes through the vision encoder, incurring substantial computational overhead. We therefore empirically examined whether sensitivity rankings obtained at a lightweight intermediate stage could serve as a reliable proxy for those derived from the full vision encoder. Intuitively, the projection layer is a natural choice since tokens here already capture high-level semantics from the vision encoder and are directly aligned with the language model. Our analysis revealed strong alignment between rankings computed from the vision encoder and those from the projection layer (Section 3.2).
@@
Motivated by this finding, we propose a training-free VLM token pruning method named ZOO-Prune (ZerOth-Order gradient estimation for token pruning). Our method measures each token’s influence by injecting Gaussian noise at a lightweight projection layer and estimating the resulting gradient norms, which we define as sensitivity. To reduce redundancy while retaining informative tokens, we introduce Sensitivity-Aware Diversity Selection, which prioritizes tokens with high sensitivity and ensures sufficient feature diversity, inspired by (Alvar et al., 2025). By jointly considering sensitivity and diversity, ZOO-Prune produces pruned token subsets that preserve task-relevant information and enable effective compression even under aggressive pruning regimes (Fig. 1).
Our main contributions are as follows:
@@
  • We propose ZOO-Prune, a training-free pruning framework that unifies sensitivity and diversity, ensuring that pruned tokens are not only highly informative but also complementary, overcoming the limitations of importance-only or diversity-only methods.
@@ 2.1 Vision-Language Models
Large Multimodal Models (LMMs), particularly Vision-Language Models (VLMs) (Liu et al., 2023; Dai et al., 2023; Chen et al., 2024b; Liu et al., 2024b; Bai et al., 2023), have demonstrated remarkable capabilities in multimodal reasoning and dialogue. Pioneering architectures like LLaVA (Liu et al., 2023) established a successful paradigm by aligning a pre-trained vision encoder (e.g., CLIP ViT (Radford et al., 2021)) with an instruction-tuned LLM through a simple projection layer. This design, further refined in subsequent works such as LLaVA family (Liu et al., 2024a; b), InternVL series (Chen et al., 2024b; Wang et al., 2024; Zhu et al., 2025; Wang et al., 2025), and Qwen-VL (Bai et al., 2023), enables strong visual understanding but introduces a significant computational challenge. These models typically encode a single image into hundreds of visual tokens, leading to substantial inference overhead. The computational burden is exacerbated when handling higher resolutions; for instance, LLaVA (Liu et al., 2023; 2024a; 2024b) typically encodes $\times336$ images into 576 tokens, and up to 2880 tokens at $\times672$ resolution. The inherent redundancy within these extensive visual token sequences has motivated research into visual token pruning as a key approach for efficient VLMs inference.
@@ 2.2 Visual Token Pruning for VLMs
Visual token pruning methods are proposed to reduce the inference complexity of large VLMs by removing redundant visual representations. Existing pruning methods can be broadly divided into two categories: those that require fine-tuning or calibration, and those that are entirely training-free. The first category relies on additional data or adaptation to guide token reduction. CrossGET (Shi et al., 2024) and MADTP (Cao et al., 2024) introduce modality-specific tokens to align cross-modal features and drive token selection, while DeCo (Yao et al., 2024) employs adaptive pooling to decouple token compression from semantic abstraction at the patch level. VTW (Lin et al., 2025) removes all vision tokens after a specific layer, identified using a small calibration set and a KL-divergence criterion. FitPrune (Ye et al., 2025) reduces visual tokens in the multi-head attention of each layer via binary search, guided by attention statistics collected from inference or calibration examples. While effective, these methods require calibration data and model-specific adaptation, limiting their flexibility across architectures and deployment scenarios.
Training-free pruning avoids retraining and offers plug-and-play acceleration. Attention-based methods estimate token importance directly from attention magnitudes. FastV (Chen et al., 2024a), LLaVA-PruMerge (Shang et al., 2024), and VisionZip (Yang et al., 2025) remove or merge tokens based on early attention maps, while SparseVLM (Zhang et al., 2025) leverages question-driven cross-attention to induce dynamic sparsity. However, attention scores are often unstable and may retain semantically redundant tokens (Alvar et al., 2025; Lin et al., 2025), which limits performance under aggressive pruning. Another direction emphasizes feature diversity. DivPrune (Alvar et al., 2025) formulates token selection as a max-min diversity problem solved via greedy farthest-point sampling, reducing redundancy and preserving robustness at high pruning ratios. Yet by treating all tokens equally, it can overlook semantically critical regions and drop task-relevant information. To overcome these limitations, we propose a zeroth-order sensitivity estimator that quantifies each token’s effect on the output without backpropagation. We further combine this signal with a diversity-based selection strategy, ensuring that pruning retains the most influential tokens while still covering complementary visual content, which leads to a stronger balance between efficiency and accuracy.
@@ 3 Methodology
3.1 Background: Zeroth-Order Gradient Estimation
Zeroth-order (ZO) optimization provides a gradient-free alternative to first-order methods by using only forward queries of a function. It is particularly useful when exact gradient computation is infeasible, such as in black-box optimization (Sawada et al., 2025), adversarial attacks (Chen et al., 2017), or efficient fine-tuning of large models (Park et al., 2025; Zhang et al., 2024b). By relying solely on forward evaluations, ZO methods circumvent the need for backpropagation, thereby reducing memory costs and enabling applications to complex or non-differentiable modules.
A widely used estimator is the randomized gradient estimator (RGE) (Duchi et al., 2015; Nesterov & Spokoiny, 2017), which approximates gradients by finite differences along random directions. Given a function $f:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$, the central-difference RGE with $m$ queries is
$\widehat{\nabla}f(x)\;=\;\frac{1}{m}\sum_{j=1}^{m}\frac{f(x+hu_{j})-f(x-hu_{j})}{2h}\,u_{j},\quad u_{j}\sim\mathcal{N}(0,I_{d}),$ (1)
@@ 3.1 Background: Zeroth-Order Gradient Estimation
3.2 Observation: Projector as a Proxy for End-to-End Vision Sensitivity
Original zeroth-order methods aim to reconstruct the full gradient vector, but in our setting, we only require a relative ranking of token importance. A naive approach would apply RGE directly to the vision encoder, incurring substantial computational overhead due to additional end-to-end forward passes. For instance, suppose an image is tokenized into $n=500$ visual tokens with $m=64$ random perturbation directions per token. Since RGE requires two forward passes per direction, the total cost scales as $2nm$ forward passes, amounting to $\sim 6.4\times 10^{6}$ GFLOPs, which is clearly prohibitive.
To reduce this cost, we empirically evaluated whether sensitivity rankings obtained at a lightweight intermediate stage (i.e., projection layer) could serve as a reliable proxy for those derived from the full vision encoder. In Fig. 2, we rank token sensitivities based on RGE and visualize the Spearman rank correlations across two datasets. Our analysis shows a strong alignment between token importance rankings computed from the vision encoder outputs and those obtained from the projection layer. Beyond empirical evidence, the projection layer can be seen as a modality-aligning bottleneck. It consolidates high-level semantic information from the vision encoder and maps it into the language embedding space, naturally emphasizing tokens that are important for downstream predictions. Since token pruning only requires relative importance rather than exact gradients, this layer provides a compact, semantically meaningful proxy that preserves token sensitivity rankings.
@@ 3.2 Observation: Projector as a Proxy for End-to-End Vision Sensitivity
3.3 ZOO-Prune: ZerOth-Order gradient estimation for token pruning
Fig. 3 illustrates the overall ZOO-Prune framework. Given the outputs of the vision encoder, our method first computes token-level sensitivity using zeroth-order gradient estimation and then uses these scores in a diversity-aware selection procedure to produce a subset of tokens that is passed to the LLM. Mathematically, let $X\in\mathbb{R}^{N_{v}\times d_{v}}$ denote $N_{v}$ vision tokens with $d_{v}$ dimension, and
$Z=M(X)\in\mathbb{R}^{N_{v}\times d_{l}}$
the projected embeddings through the multimodal projection layer $M:\mathbb{R}^{d_{v}}\to\mathbb{R}^{d_{l}}$. We sample $m$ random perturbation directions $\{u_{j}\}_{j=1}^{m}$, where $u_{j}\sim\mathcal{N}(0,I_{d_{v}})$ normalized to unit norm. For each token $i$, we measure the symmetric finite-difference response:
@@ 3.3 ZOO-Prune: ZerOth-Order gradient estimation for token pruning
where $x_{i}\in\mathbb{R}^{d_{v}}$ is the $i$-th vision token. We then define the token sensitivity of token $i$ as the average response magnitude:
@@
Let $M:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be differentiable at $x\in\mathbb{R}^{n}$ with Jacobian $J(x)=\nabla M(x)$. Let $u\sim\mathcal{N}(0,I_{n})$ be an isotropic Gaussian perturbation and $h>0$ a small step size. Define the finite-difference sensitivity $S(x)=\mathbb{E}_{u}\left[\left\|\frac{M(x+hu)-M(x-hu)}{2h}\right\|_{2}\right].$ Then, for sufficiently small $h$,
$S(x)=\mathbb{E}_{u}\big[\|J(x)u\|_{2}\big]+O(h^{2}).$
The detailed proof is provided in Appendix B. The proposition shows that, unlike traditional RGE, which estimates gradient direction, our sensitivity metric captures the magnitude of local response. Specifically, $\mathbb{E}_{u}[\|J(x)u\|_{2}]$ quantifies how much the output changes on average under random perturbations, providing a scalar measure of token influence.
Sensitivity-aware diversity selection. While sensitivity captures the most informative tokens, it does not by itself enforce coverage over diverse visual content. To reduce redundancy, we integrate a diversity criterion inspired by DivPrune (Alvar et al., 2025). Let $Z_{i}$ denote the vision feature of token $i$, and $\mathcal{P}$ the set of already selected tokens. We define
Algorithm 1 ZOO-Prune: ZerOth-Order Sensitivity-Aware Token Pruning
@@
3:% —— [ZOO-based Sensitivity Estimation] ———-
@@
9:% —— [Sensitivity-Aware Diversity Selection ] ——

### v2

# ZOO-Prune: Training-Free Token Pruning via Zeroth-Order Gradient Estimation in Vision-Language Models

@@ Abstract
Figure 1: Illustration of training-free VLM token pruning methods. (a) Attention-based methods select tokens using attention scores, but often retain redundant tokens. (b) Diversity-based methods select tokens with different features to maximize coverage but may lose tokens located in semantically relevant regions (e.g., around the monitor, highlighted in yellow). (c) Our method employs zeroth-order gradient estimation to quantify token sensitivity and integrates these scores into a diversity objective. (d) Accuracy comparison with LLaVA-NeXT-7B across 9 benchmarks, showing that ours outperforms both VisionZip (attention-based) and DivPrune (diversity-based).
@@ 1 Introduction
Large Vision–Language Models (VLMs) [3, 25, 51] achieve strong multimodal understanding but at a substantial computational cost. A major contributor to this cost is the large number of visual tokens produced by modern vision backbones. For example, the vision encoder in LLaVA-1.5 [24] generates up to 576 tokens for a single image, whereas the text side often contains only a few tokens, such as “Describe this image in a short sentence”. This imbalance leads to high inference latency and memory overhead [10, 1]. To address this issue, token pruning has emerged as a practical solution that selectively removes less informative visual tokens at inference time. Prior work has shown that aggressive token reduction can yield large speedups with only modest drops in accuracy [49, 42, 2].
@@
Recent research on VLM token pruning has increasingly focused on training-free schemes. Unlike works that require calibration data or fine-tuning [19, 21, 23], training-free approaches prune tokens directly at inference. These methods can be broadly categorized into two groups. (1) Attention-based approaches score tokens using attention magnitudes; however, attention often concentrates on background regions [16] and tends to retain redundant tokens with overlapping content [42, 34]. For example, in an image of a laptop on a desk (Fig. 1), attention-based pruning may keep many redundant background tokens while missing tokens near the monitor that are important for answering questions. (2) Diversity-based approaches [1, 12, 31] select tokens by maximizing feature diversity, measuring pairwise distances between token embeddings. However, because these methods prioritize diversity uniformly across all tokens without explicitly considering task-relevant cues, they may discard tokens from visually salient regions. In the same laptop example, diversity-based pruning may select feature-diverse tokens, but not necessarily those near the monitor, which are crucial for reasoning about nearby objects.
Given these limitations, selecting a smaller subset of tokens while preserving visual information without degrading task performance remains a challenging problem. To tackle this, we investigate a novel metric for scoring each token, explicitly considering token sensitivity (i.e., the degree to which small perturbations to a token change the model’s output). Prior work has shown that attention weights do not necessarily correlate with a token’s actual impact on the model output [6, 40], motivating the use of sensitivity-based measures that directly quantify output changes under token perturbations. [9]. Unfortunately, directly computing token sensitivities via gradients is costly. It also requires ground-truth outputs to define a supervised loss, which is unavailable during inference-time pruning. These factors motivate us to use zeroth-order gradient estimation [29], which quantifies how perturbations to input tokens affect the model output using only forward passes, avoiding backpropagation.
@@
However, a naive application of zeroth-order estimation would require additional forward passes through the vision encoder, incurring substantial computational overhead. We therefore empirically examined whether sensitivity rankings obtained at a lightweight intermediate stage could serve as a reliable proxy for those derived from the full vision encoder. Intuitively, the projection layer is a natural choice since tokens here already capture high-level semantics from the vision encoder and are directly aligned with the language model [8, 37]. Our analysis revealed strong alignment between rankings computed from the vision encoder and those from the projection layer (Section 3.2).
@@
Motivated by this finding, we propose a training-free VLM token pruning method named ZOO-Prune (ZerOth-Order gradient estimation for token pruning). Our method measures each token’s influence by injecting Gaussian noise at a lightweight projection layer and estimating the resulting gradient norms, which we define as sensitivity. To reduce redundancy while retaining informative tokens, we introduce sensitivity-aware diversity selection, which prioritizes tokens with high sensitivity and ensures sufficient feature diversity, inspired by [1]. By jointly considering sensitivity and diversity, ZOO-Prune produces pruned token subsets that preserve task-relevant information and enable effective compression even under aggressive pruning regimes (Fig. 1). Our main contributions are as follows:
@@
  • We propose ZOO-Prune, a training-free pruning framework that unifies sensitivity and diversity, ensuring that pruned tokens are not only highly informative but also complementary, overcoming the limitations of attention-only or diversity-only methods.
@@ 2.1 Vision-Language Models
Large Multimodal Models (LMMs), particularly Vision-Language Models (VLMs) [26, 15, 13, 25, 3], have demonstrated remarkable capabilities in multimodal reasoning and dialogue. Pioneering architectures like LLaVA [26] established a successful paradigm by aligning a pre-trained vision encoder (e.g., CLIP ViT [32]) with an instruction-tuned LLM through a simple projection layer. This design, further refined in subsequent works such as LLaVA family [24, 25], InternVL series [13, 38, 51, 39], and Qwen-VL [3], enables strong visual understanding but introduces a significant computational challenge. These models typically encode a single image into hundreds of visual tokens, leading to substantial inference overhead. The computational burden is exacerbated when handling higher resolutions; for instance, LLaVA [26, 24, 25] typically encodes $\times336$ images into 576 tokens, and up to 2880 tokens at $\times672$ resolution. The inherent redundancy within these extensive visual token sequences has motivated research into visual token pruning as a key approach for efficient VLMs inference.
@@ 2.2 Visual Token Pruning for VLMs
Visual token pruning methods are proposed to reduce the inference complexity of large VLMs by removing redundant visual representations [19, 21, 23]. Existing methods can be broadly divided into two categories: those that require fine-tuning or calibration, and those that are entirely training-free. The first category relies on additional data or adaptation to guide token reduction. CrossGET [35] and MADTP [7] introduce modality-specific tokens to align cross-modal features and drive token selection, while DeCo [43] employs adaptive pooling to decouple token compression from semantic abstraction at the patch level. VTW [23] removes all vision tokens after a specific layer, identified using a small calibration set and a KL-divergence criterion. FitPrune [44] reduces visual tokens in the multi-head attention of each layer via binary search, guided by attention statistics collected from inference or calibration examples. While effective, these methods require calibration data and model-specific adaptation, limiting their flexibility across architectures and deployment scenarios.
Training-free pruning avoids retraining and offers plug-and-play acceleration. Attention-based methods stimate token importance directly from attention magnitudes. FastV [10], LLaVA-PruMerge [34], ToMe [5] and VisionZip [42] remove or merge tokens based on early attention maps, while SparseVLM [49] and PyramidDrop [41] leverage question-driven cross-attention to induce dynamic sparsity. However, attention scores are often unstable and may retain semantically redundant tokens [1, 23], which limits performance under aggressive pruning. Another direction emphasizes feature diversity. DivPrune [1] formulates token selection as a max-min diversity problem solved via greedy farthest-point sampling, reducing redundancy and preserving robustness at high pruning ratios. Yet by treating all tokens equally, it can overlook semantically critical regions and drop task-relevant information. To overcome these limitations, we propose a zeroth-order sensitivity estimator that quantifies each token’s effect on the output without backpropagation. We further combine this sensitivity measure with a diversity-based selection strategy so that pruning preserves the most influential tokens while still capturing complementary visual content.
@@ 3 Methodology
3.1 Zeroth-Order Gradient Estimation
Zeroth-order (ZO) optimization offers a gradient-free alternative to first-order methods by using only forward queries of a function. It is particularly useful when exact gradient computation is infeasible, such as in black-box optimization [33], adversarial attacks [11], or efficient fine-tuning of large models [30, 48]. In our setting, computing exact gradients would require full backpropagation through a large-scale LLM, which is prohibitively expensive at inference time. By relying solely on forward evaluations, ZO methods circumvent the need for backpropagation, thereby reducing memory costs and enabling applications to complex or non-differentiable modules.
A widely used estimator is the randomized gradient estimator (RGE) [17, 29], which approximates gradients by finite differences along random directions. Given a function $f:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$, the central-difference RGE with $m$ queries is:
$\widehat{\nabla}f(x)\;=\;\frac{1}{m}\sum_{j=1}^{m}\frac{f(x+hu_{j})-f(x-hu_{j})}{2h}\,u_{j},$ (1)
$\quad u_{j}\sim\mathcal{N}(0,I_{d}),$ (2)
@@ 3.2 Projector as a Proxy for Visual Sensitivity
Original zeroth-order methods aim to reconstruct the full gradient vector, but in our setting, we only require a relative ranking of token importance. A naive approach would apply RGE directly to the vision encoder, incurring substantial computational overhead due to the additional end-to-end forward passes. For instance, suppose an image is tokenized into $n=500$ visual tokens with $m=64$ random perturbation directions per token. Since RGE requires two forward passes per direction, the total cost scales as $2nm$ forward passes, amounting to $\sim 6.4\times 10^{6}$ GFLOPs, which is clearly prohibitive.
To reduce this cost, we empirically evaluated whether sensitivity rankings obtained at a lightweight intermediate stage (i.e., projection layer) could serve as a reliable proxy for those derived from the full vision encoder. In Fig. 2, we rank token sensitivities based on RGE and visualize the Spearman rank correlations across two datasets. Our analysis shows a feasible alignment between token importance rankings computed from the vision encoder outputs and those obtained from the projection layer. Beyond empirical evidence, the projection layer can be seen as a modality-aligning bottleneck. It consolidates high-level semantic information from the vision encoder and maps it into the language embedding space, naturally emphasizing tokens that are important for downstream predictions. Since token pruning only requires relative importance rather than exact gradients, this layer provides a compact, semantically meaningful proxy that preserves token sensitivity rankings.
@@
3.3 ZOO-Prune
Fig. 3 illustrates the overall ZOO-Prune (ZerOth-Order gradient estimation for token pruning) framework. Given the outputs of the vision encoder, our method first computes token-level sensitivity using zeroth-order gradient estimation and then uses these scores in a diversity-aware selection procedure to produce a subset of tokens that is passed to the LLM. Mathematically, let $X\in\mathbb{R}^{N_{v}\times d_{v}}$ denote $N_{v}$ vision tokens with $d_{v}$ dimension, and $Z=M(X)\in\mathbb{R}^{N_{v}\times d_{l}}$ the projected embeddings through the multimodal projection layer $M:\mathbb{R}^{d_{v}}\to\mathbb{R}^{d_{l}}$. We sample $m$ random perturbation directions $\{u_{j}\}_{j=1}^{m}$, where $u_{j}\sim\mathcal{N}(0,I_{d_{v}})$ normalized to unit norm. For each token $i$, we measure the symmetric finite-difference response:
@@ 3.3 ZOO-Prune
where $x_{i}\in\mathbb{R}^{d_{v}}$ is the $i$-th vision token. We define the token sensitivity of token $i$ as the average response magnitude:
@@
Let $M:\!\mathbb{R}^{n}\!\to\!\mathbb{R}^{m}$ be differentiable at $x\!\in\!\mathbb{R}^{n}$ with Jacobian $J(x)\!=\!\nabla M(x)$. Let $u\sim\mathcal{N}(0,I_{n})$ be an isotropic Gaussian perturbation and h>0h\!>\!0 a small step size. Define the finite-difference sensitivity $S(x)\!=\!\mathbb{E}_{u}\left[\left\|\frac{M(x+hu)-M(x-hu)}{2h}\right\|_{2}\right].$ Then, for sufficiently small $h$:
$S(x)=\mathbb{E}_{u}\big[\|J(x)u\|_{2}\big]+O(h^{2}).$ (5)
The detailed proof is provided in Appendix B. The proposition shows that, unlike traditional RGE, which estimates gradient direction, our sensitivity metric captures the magnitude of local response. Specifically, $\mathbb{E}_{u}[\|J(x)u\|_{2}]$ quantifies how much the output changes on average under random perturbations, providing a scalar measure of the token influence.
Algorithm 1 ZOO-Prune
@@
3:% ———– [ZOO-based Sensitivity Estimation] ————-
@@
9:% ———– [Sensitivity-Aware Diversity Selection ] ———
@@
Sensitivity-aware diversity selection.
While sensitivity captures the most informative tokens, it does not by itself enforce coverage over diverse visual content. To reduce redundancy, we integrate a diversity criterion inspired by DivPrune [1]. Let $Z_{i}$ denote the vision feature of token $i$, and $\mathcal{P}$ the set of already selected tokens. We define
4 Experiments
[…]